# Quasi Zigzag Persistence: A Topological Framework for Analyzing Time-Varying Data

### v1

@@ 6 Conclusion
In this paper, we proposed QZPH as a framework to capture both static and dynamic topological features in time-varying data. We proposed Zz-Gril, a stable and computationally efficient topological invariant to address the challenges of integrating MPH and ZPH. Through applications in various domains, including sleep-stage detection, we showed that augmenting machine learning models with Zz-Gril improves the performance. These results highlight the potential of integrating topological information to address complex challenges while analyzing time-evolving data. The differentiability of Zz-Gril is an interesting research avenue. Further, it would also be interesting to study how other multiparameter methods adapt to the QZPH framework.
@@ 7 Acknowledgement
Appendix A Formal definitions and proofs
@@
The interleaving distance on $\mathbb{R}^{n}$-indexed persistence modules is an extension of the same distance defined for $\mathbb{R}$-indexed persistence modules Chazal et al. (2009). We briefly recall the definition here. We refer the readers to Botnan & Lesnick (2018) for additional details.
@@
The $u$-shift functor $(-)_{u}\colon\mathbf{vec}^{\mathbb{R}^{n}}\to\mathbf{vec}^{\mathbb{R}^{n}}$, for $u\in\mathbb{R}^{n}$, is defined as follows:
@@
    Let $M,N\in\mathbf{vec}^{\mathbb{R}^{n}}$. Let $F\colon M\to N$ be a morphism. Then, the corresponding morphism $F_{u}\colon M_{u}\to N_{u}$ is defnied as $F_{u}(x)=F(x+u)\colon M_{u}(x)\to N_{u}(x)$ for all $x\in\mathbb{R}^{n}$.
@@
Let $M,N\in\mathbf{vec}^{\mathbb{R}^{n}}$. Let $\epsilon\in[0,\infty)$ be given. We will denote $(-)_{\bm{\epsilon}}$ to be the shift functor corresponding to the vector $\bm{\epsilon}=\epsilon(1,1,\ldots,1)$. We say $M$ and $N$ are $\epsilon$-interleaved if there are natural transformations $F\colon M\to N_{\bm{\epsilon}}$ and $G\colon N\to M_{\bm{\epsilon}}$ such that
@@
where $\varphi_{M}^{u}\colon M\to M_{u}$ is the natural transformation whose restriction to each $M(x)$ is the linear map $M(x\leq x+u)$ for all $x\in\mathbb{R}^{n}$.
@@
Let $P$ and $Q$ be two posets. Let $F\colon P\to Q$ be a functor. Let $P[F\leq q]$ denote the set $P[F\leq q]\coloneqq\{p\in P\colon F(p)\leq q\}$. Given a persistence module $M\colon P\to\mathbf{vec}$, the left Kan extension of $M$ along $F$ is a functor $\text{Lan}_{F}(M)\colon Q\to\mathbf{vec}$ given by
@@
Let $\mathbb{U}\subset\mathbb{R}^{\mathsf{op}}\times\mathbb{R}$ denote the poset $\mathbb{U}\coloneqq\{(a,b)\colon a\leq b\}$. Let $\mathfrak{i}\colon\mathbb{ZZ}\to\mathbb{R}^{\mathsf{op}}\times\mathbb{R}$ denote the inclusion. Let $(-)|_{\mathbb{U}}\colon\mathbf{vec}^{\mathbb{R}^{\mathsf{op}}\times\mathbb{R}}\to\mathbf{vec}^{\mathbb{U}}$ denote the restriction. Then, the block extension functor $E\colon\mathbf{vec}^{\mathbb{ZZ}}\to\mathbf{vec}^{\mathbb{U}}$ is defined as
@@
We extend this definition of interleaving distance to quasi zigzag persistence modules. Let $\mathbb{UU}\subset\mathbb{R}^{\mathsf{op}}\times\mathbb{R}\times\mathbb{R}$ be the analog of $\mathbb{U}$, i.e, $\mathbb{UU}\coloneqq\{(a,b,c)\colon a\leq b\}$. Let $\tilde{\mathfrak{i}}\colon\mathbb{ZZ}\times\mathbb{Z}\to\mathbb{R}^{\mathsf{op}}\times\mathbb{R}\times\mathbb{R}$ denote the inclusion. Then, define $\tilde{E}\coloneqq(-)|_{\mathbb{UU}}\circ\text{Lan}_{\tilde{\mathfrak{i}}}(\circ)$, analogous to the block extension functor.
@@
Let $(\mathbb{ZZ}\times\mathbb{Z})$ denote the collection of all subposets in $\mathbb{ZZ}\times\mathbb{Z}$ such that their corresponding subposets in $\mathbb{Z}^{2}$ are intervals. Let $M$ and $N$ be two quasi zigzag persistence modules. The erosion distance is defined as:
$\begin{split}d_{\mathcal{E}}(M,N)\coloneqq\inf\limits_{\epsilon\geq 0}\{&\forall I\in\textbf{I}(\mathbb{ZZ}\times\mathbb{Z}),\\ &\mathsf{rk}^{M}(I)\geq\mathsf{rk}^{N}(I^{\epsilon})\text{ and }\\ &\mathsf{rk}^{N}(I)\geq\mathsf{rk}^{M}(I^{\epsilon})\}\end{split}$
Definition A.11.
Let $\mathcal{L}$ denote the collection of all worms in $\mathbb{ZZ}\times\mathbb{Z}$. Let $M$ and $N$ be quasi zigzag persistence modules. The erosion distance can be defined as:
$\begin{split}d_{\mathcal{E}}^{\mathcal{L}}(M,N)\coloneqq\inf\limits_{\epsilon\geq 0}\{&\forall\boxed{\mathbf{p}}^{2}_{\delta}\in\textbf{I}(\mathbb{ZZ}\times\mathbb{Z}),\\ &\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta}\right)\geq\mathsf{rk}^{N}\left(\boxed{\mathbf{p}}^{2}_{\delta+\epsilon}\right)\text{ and }\\ &\mathsf{rk}^{N}\left(\boxed{\mathbf{p}}^{2}_{\delta}\right)\geq\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta+\epsilon}\right)\}.\end{split}$

### v2

@@ 6 Conclusion
In this paper, we proposed QZPH as a framework to capture both static and dynamic topological features in time-varying data. We proposed Zz-Gril, a stable and computationally efficient topological invariant to address the challenges of integrating MPH and ZPH. Through applications in various domains, including sleep-stage detection, we showed that augmenting machine learning models with Zz-Gril improves the performance. These results highlight the potential of integrating topological information to address complex challenges while analyzing time-evolving data.
Limitations. One of the limitations of this method, like most other persistent homology based methods, is its scalability. With the increasing size of the bifiltrations, the number of sampled center points to compute Zz-Gril needs to increase. Consequently, the number of zigzag persistence computations and the length of the zigzag filtrations over which they need to be performed increase. However, recent advances in persistent homology computations in the context of ML Nigmetov & Morozov (2024) give us the hope.
@@ 7 Acknowledgement
Appendix A Missing Definitions and Proofs
@@
The interleaving distance on $\mathbb{R}^{n}$-indexed persistence modules is known and well-defined. We briefly recall the definition here. We refer the readers to Botnan & Lesnick (2018) for additional details.
@@
The $u$-shift functor $(-)_{u}:\mathbf{vec}^{\mathbb{R}^{n}}\to\mathbf{vec}^{\mathbb{R}^{n}}$, for $u\in\mathbb{R}^{n}$, is defined as follows:
@@
    Let $M,N\in\mathbf{vec}^{\mathbb{R}^{n}}$. Let $F:M\to N$ be a morphism. Then, the corresponding morphism $F_{u}:M_{u}\to N_{u}$ is defnied as $F_{u}(x)=F(x+u):M_{u}(x)\to N_{u}(x)$ for all $x\in\mathbb{R}^{n}$.
@@
Let $M,N\in\mathbf{vec}^{\mathbb{R}^{n}}$. Let $\epsilon\in[0,\infty)$ be given. We will denote $(-)_{\bm{\epsilon}}$ to be the shift functor corresponding to the vector $\bm{\epsilon}=\epsilon(1,1,\ldots,1)$. We say $M$ and $N$ are $\epsilon$-interleaved if there are natural transformations $F:M\to N_{\bm{\epsilon}}$ and $G:N\to M_{\bm{\epsilon}}$ such that
@@
where $\varphi_{M}^{u}:M\to M_{u}$ is the natural transformation whose restriction to each $M(x)$ is the linear map $M(x\leq x+u)$ for all $x\in\mathbb{R}^{n}$.
@@
Let $P$ and $Q$ be two posets. Let $F:P\to Q$ be a functor. Let $P[F\leq q]$ denote the set $P[F\leq q]\coloneqq\{p\in P:F(p)\leq q\}$. Given a persistence module $M:P\to\mathbf{vec}$, the left Kan extension of $M$ along $F$ is a functor $\text{Lan}_{F}(M):Q\to\mathbf{vec}$ given by
@@
Let $\mathbb{U}\subset\mathbb{R}^{\mathsf{op}}\times\mathbb{R}$ denote the poset $\mathbb{U}\coloneqq\{(a,b):a\leq b\}$. Let $\mathfrak{i}:\mathbb{ZZ}\to\mathbb{R}^{\mathsf{op}}\times\mathbb{R}$ denote the inclusion. Let $(-)|_{\mathbb{U}}:\mathbf{vec}^{\mathbb{R}^{\mathsf{op}}\times\mathbb{R}}\to\mathbf{vec}^{\mathbb{U}}$ denote the restriction. Then, the block extension functor $E:\mathbf{vec}^{\mathbb{ZZ}}\to\mathbf{vec}^{\mathbb{U}}$ is defined as
@@
We extend this definition of interleaving distance to quasi zigzag persistence modules. Let $\mathbb{UU}\subset\mathbb{R}^{\mathsf{op}}\times\mathbb{R}\times\mathbb{R}$ be the analog of $\mathbb{U}$, i.e, $\mathbb{UU}\coloneqq\{(a,b,c):a\leq b\}$. Let $\tilde{\mathfrak{i}}:\mathbb{ZZ}\times\mathbb{Z}\to\mathbb{R}^{\mathsf{op}}\times\mathbb{R}\times\mathbb{R}$ denote the inclusion. Then, define $\tilde{E}\coloneqq(-)|_{\mathbb{UU}}\circ\text{Lan}_{\tilde{\mathfrak{i}}}(\circ)$, analogous to the block extension functor.
@@
Let $I$ denote a subposet in $\mathbb{ZZ}\times\mathbb{Z}$. For ϵ+italic-$\epsilon\in\mathbb{Z}_{+}$, the $\epsilon$-thickening of $I$ is defined as
$I^{\epsilon}\coloneqq\left\{\mathbf{r}\in\mathbb{ZZ}\times\mathbb{Z}:\exists\mathbf{q}\in I\text{ such that }||\mathbf{r}-\mathbf{q}||_{\infty}\leq\epsilon\right\}.$
Recall that $(\mathbb{ZZ}\times\mathbb{Z})$ denotes the collection of all subposets in $\mathbb{ZZ}\times\mathbb{Z}$ such that their corresponding subposets in $\mathbb{Z}^{2}$ are intervals.
@@
Let $M$ and $N$ be two quasi zigzag persistence modules. The erosion distance is defined as:
$d_{\mathcal{E}}(M,N)\coloneqq\inf\{\epsilon\geq 0:\forall I\in\textbf{I}(\mathbb{ZZ}\times\mathbb{Z}),\mathsf{rk}^{M}(I)\geq\mathsf{rk}^{N}(I^{\epsilon})\text{ and }\mathsf{rk}^{N}(I)\geq\mathsf{rk}^{M}(I^{\epsilon})\}.$